PERIODIC SOLUTIONS OF A SUBQUADRATIC NONCONVEX SECOND ORDER HAMILTONIAN SYSTEM by Lavoslav Čaklović, Zagreb, Croatia

Abstract. The second order hamiltonian system is considered with T-periodic potential that is subquadratic at infinity. The existence of T-periodic solution is proved using variational methods. Mathematics subject classification: 34L30, 58E05, 58F22. Keywords: Hamiltonian system, PS-condition, deformation lemma.